# Multigranular Evaluation for Brain Visual Decoding

### v1

@@ 3.3 BASIC-L
NSD allen2022massive
SDRecon takagi2023improving 9.03 13.06 38.90 49.97 15.08 21.43 17.84 1.07 4.38 6.79 11.81
BrainDiffuser ozcelik2023brain 17.96 20.21 38.98 45.85 18.66 20.78 20.09 1.94 7.86 7.82 16.65
MindEye scotti2023reconstructing 19.20 22.79 45.53 55.17 18.73 21.94 20.36 1.99 7.49 8.26 17.03
DREAM xia2024dream 23.62 26.10 46.03 57.10 21.15 24.13 21.41 2.32 9.22 8.79 19.57
MindEye2 scotti2024mindeye2 25.29 26.27 47.93 57.52 24.33 25.68 24.09 3.45 12.32 11.03 22.16
MindBridge wang2024mindbridge 16.24 19.04 40.51 48.95 16.87 19.81 18.61 1.54 6.30 6.78 15.00
UMBRAE xia2024umbrae 21.33 24.62 40.96 48.53 18.94 21.16 20.29 2.15 8.43 8.47 17.89
NeuroPictor huo2024neuropictor 29.45 31.29 47.79 56.48 27.97 29.57 26.47 4.08 17.17 15.84 25.88
NeuroVLA shen2024neuro 16.03 20.85 44.09 54.42 13.84 17.54 17.57 1.34 4.38 5.57 13.54
SepBrain wang2024unibrain 21.22 23.72 45.06 53.47 20.88 23.32 21.98 2.51 8.79 8.98 18.84
UniBrain wang2024unibrain 15.24 18.02 37.48 45.09 14.99 17.75 17.45 1.06 5.54 6.07 13.79
STTM liu2025see 27.31 29.44 49.35 59.05 24.61 26.37 24.19 3.50 12.55 11.65 22.90
MindTuner gong2025mindtuner 15.38 16.66 40.74 49.02 21.46 24.50 21.49 1.97 8.16 8.13 16.98
BrainGuard tian2025brainguard 25.90 28.07 49.22 58.99 23.34 25.21 23.98 3.29 10.82 10.32 21.76
EEG-Things  grootswagers2022human
ATM li2024visual 13.87 18.77 39.46 50.14 22.85 34.69 22.02 2.15 11.13 17.07 17.60
CognitionCapturer zhang2025cognitioncapturer 10.26 12.71 33.31 40.67 21.37 29.84 20.75 1.37 13.08 17.26 16.22
CC2017 wen2018neural
MinD-Video chen2023cinematic 12.82 24.20 44.89 53.50 20.29 33.57 19.87 3.53 6.83 14.69 15.25
NeuroClips gong2024neuroclips 24.37 31.59 65.51 73.39 28.23 36.00 28.74 7.73 9.82 13.14 23.52
DecoFuse li2025decofuse 12.32 18.07 49.87 59.46 22.33 34.63 20.87 2.74 4.06 9.80 15.31
SEED-DV liu2024eeg2video
EEG2Video liu2024eeg2video 27.77 32.82 57.26 69.97 22.93 31.96 23.41 3.51 3.10 7.10 20.54
fMRI-Shape gao2024mind3d
MinD-3D gao2024mind3d 5.69 8.10 5.69 8.10 19.91 29.35 19.38 1.45 15.96 25.30 14.72
MinD-3D++ gao2024mind3d++ 3.80 5.23 3.80 5.23 16.03 35.83 17.96 1.83 14.94 36.19 12.62
EEG-3D guo2025neuro3d
Neuro-3D guo2025neuro3d 2.39 3.65 2.39 3.65 13.77 25.92 12.25 1.02 12.33 18.56 9.69
4 Experiments
This section presents an experimental evaluation of existing brain visual decoding methods across multiple benchmark datasets. Experimental Details (Sec. 4.1) introduces the datasets, decoding methods, implementation settings, and additional evaluation protocols, while Experimental Results (Sec. 4.2) reports the quantitative performance using our proposed evaluation framework.
4.1 Experimental Setup
Stimulus-neuroimaging datasets and methods.
We benchmark brain visual decoding methods across six representative datasets spanning three major visual modalities (images, videos, and 3D shapes ) and two neuroimaging techniques (fMRI and EEG). These datasets include: the fMRI-image dataset NSD, which has been used in various decoding methods for visual image reconstruction takagi2023improving ; ozcelik2023brain ; scotti2023reconstructing ; scotti2024mindeye2 ; xia2024dream ; xia2024umbrae ; shen2024neuro ; the fMRI-video dataset CC2017 wen2018neural , which has been employed in methods exploring dynamic stimuli decoding li2025decofuse ; chen2023cinematic ; huo2024neuropictor ; and the fMRI-3D dataset MinD-3D gao2024mind3d , focusing on shape-level decoding from object stimuli, with related methods examining brain-based shape reconstruction gao2024mind3d ; gao2024mind3d++ . EEG-based datasets such as EEG-Things grootswagers2022human , SEED-DV liu2024eeg2video , and EEG-3D guo2025neuro3d have been used to decode brain activities for images grootswagers2022human ; li2025decofuse ; chen2023cinematic ; huo2024neuropictor , videos liu2024eeg2video , and 3D shapes guo2025neuro3d . For each dataset, we evaluate all available brain visual decoding methods , as acknowledged, with open-source implementations or author-shared reconstruction results within our unified evaluation framework. These methods represent the state of the art in brain visual decoding and reflect recent methodological trends using pretrained generative models radford2021learning ; rombach2022high . This coverage enables standardized and multigranular comparisons across the current landscape of brain visual decoding research. Further details on datasets and methods are in Appendix A and Appendix B.
@@ 4.1 Experimental Setup
LLaVA-1.6-13B liu2023visual is set as the default MLLM for detailed description generation and semantic instance categorization. Rather than using GPT-4o or other larger but more capable models, we opt LLaVA-1.6-13B as a practical trade-off between captioning accuracy and computational efficiency. Ground truth captions are obtained from human experts through error correction, missing element addition, and hallucination removal, based on GPT-4o-generated captions, for better alignment with human judgment. For multigranular segmentation, we employ Grounded-SAM2 ren2024grounded , an open-source referring expression comprehension method, to obtain foreground, semantic, instance, and part-level segmentation masks. The default weights $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ for BASIC-H, and $\beta_{1}$, $\beta_{2}$, $\beta_{3}$ and $\beta_{4}$ for BASIC-L, are set to 4:4:2 and 3:2.5:2.5:2, respectively, reflecting the empirically informed importance of each sub-indicator in capturing semantic and structural alignment. Experiments are conducted on an NVIDIA A100 GPU. To ensure broad applicability and methodological consistency, we refrain from using MLLMs or segmentation methods specifically designed for video or 3D data. Instead, we retain an image-based framework with automatic selection of representative video frames and rendered 3D views that best capture key semantics and structure from decoded reconstructions and corresponding stimuli guo2025neuro3d ; gao2024mind3d ; liu2024eeg2video ; wen2018neural .
4.2 Experimental Results
This section presents experimental results showcasing the effectiveness of our proposed framework for brain visual decoding evaluation. The dual-perspective framework enables fine-grained and holistic comparisons, revealing method capabilities, limitations, and future development implications.
Quantitative evaluation for fMRI-to-image reconstruction on NSD allen2022massive following standard NSD and BASIC metrics is in Tab. S2. As shown, while NSD metrics saturates in scores and fails to distinguish method performance, our metrics provide clearer differentiation and more discriminative information across objects, attributes, and relations for high-level semantics, as well as salient, semantic, instance, and part granularity for low-level structure. Details for each dimension are in Tab. 2 and Tab. 3.
@@ 4.2 Experimental Results
Tab. 2 presents a comparison of object, attribute, and relation prediction performance across several datasets. While prior metrics often produce saturated or undifferentiated scores across models (see Appendix B), our proposed multigranular evaluation reveals meaningful variations in object coherence, attribute accuracy, and relational plausibility, offering a more fine-grained diagnostic lens. Here, for methods on NSD allen2022massive , NeuroVLA shen2024neuro , DREAM xia2024dream , and STTM liu2025see achieve the highest BASIC-H scores (47.88, 46.37, and 45.88, respectively), reflecting their strength in modeling rich visual semantics. This strength likely stems from detailed caption generation, complex visual semantics modeling, cross-subject training, and the combination of these factors. A closer comparison between NeuroVLA and DREAM shows that while both models reconstruct the correct instances (achieving high precision) across the three semantic dimensions, NeuroVLA misses fewer relevant instances (better recall). In contrast, SDRecon takagi2023improving , BrainDiffuser ozcelik2023brain , and UniBrain wang2024unibrain lag behind with BASIC-H scores of 35.31, 39.71, and 39.89, respectively. These methods struggle notably with the identification of object categories, attributes, and relations, indicating limitations in handling semantics or generalizing across visual categories.
@@
Tab. 3 reports performance across salient, binary, semantic, instance, and part segmentation matching scores. For decoding methods on NSD allen2022massive , NeuroPictor huo2024neuropictor leads with the highest BASIC-L score of 25.88, driven by its superior performance in instance and part segmentation. This reflects its ability to reconstruct fine-grained visual details and delineate object boundaries from neural signals. STTM liu2025see and MindEye2 scotti2024mindeye2 also perform competitively, with BASIC-L scores of 22.90 and 22.16, respectively. SDRecon takagi2023improving and NeuroVLA shen2024neuro , with BASIC-L scores of 11.81 and 13.54, respectively, show relatively poor spatial structural fidelity. BrainGuard tian2025brainguard achieves competitive scores in salient and semantic categories but performs less satisfactorily in instance and part segmentation. This indicates that while the model captures the overall structure of salient objects and categories, it faces challenges in distinguishing between multiple instances within the same category and identifying sub-component relationships.
Figure 2: BASIC performance.
Fig. 2 provides scatter plot showing the performance of visual decoding models on the NSD dataset in terms of structural alignment (BASIC-L) and semantic alignment (BASIC-H). Each point represents a model, revealing trade-offs between spatial fidelity and semantic coherence. Models closer to the top-right demonstrate stronger performance across both dimensions. For structural alignment, NeuroPictor huo2024neuropictor , STTM liu2025see , and MindEye2 scotti2024mindeye2 achieve the highest BASIC-L scores, reflecting better spatial reconstruction. In contrast, semantic alignment (BASIC-H) is best captured by NeuroVLA shen2024neuro , DREAM xia2024dream , and STTM liu2025see , indicating superior semantic preservation. More qualitative and quantitative results and analyses are provided in Appendix D and Appendix E.
4.3 Discussion
@@
Open: Our evaluation pipeline is designed to be model-agnostic, avoiding reliance on proprietary or task-specific components to ensure broad applicability and reproducibility. While models like GPT-4o are shown to have stronger multimodal performance in captioning benchmarks lu2024benchmarking ; dong2024benchmarking , they are not suitable for open, large-scale benchmarking due to API restrictions and cost. We use LLaVA-1.6-13B, which balances captioning accuracy and computational efficiency. Stable: The methods under our metrics demonstrate stable and consistent performance despite variations in (a) MLLMs, (b) prompting strategies, and (c) thresholds for box and text, as shown in Fig. 3. The relative ranking and discriminative power across decoding methods remain consistent, although absolute scores may vary slightly. The stable BASIC-H results across different MLLMs and prompts indicate that our metric reliably captures semantic performance differences among methods. Similarly, the BASIC-L scores provide consistent structural evaluation across varying text and box threshold settings, where only the highest similarity boxes exceeding the box threshold and words with similarity scores above the text threshold are considered predicted labels. Versatile: For video and 3D data evaluation, instead of using modality-specific tools, we maintain the same image-based pipeline, with automatic selection of representative video frames and 3D-rendered views to ensure compatibility across formats without compromising evaluation quality.
Figure 3: BASIC demonstrates stable and consistent performance in method evaluation across variations in (a) MLLMs, (b) prompting strategies, and (c) thresholds for box and text.
@@
This multigranular, interpretable feedback provides fine-grained diagnostic insights into brain-based visual decoding and can help uncover blind spots. Decoding methods that rely on pretrained models benefit from powerful generative capabilities but are also susceptible to systematic biases, often introducing hallucinated details such as prototypical co-occurrence patterns. For instance, savannah may be infered when the decoded scene label is giraffe even though the original stimulus was a zoo. These hallucinations tend to affect high-level contextual information—such as scene types or typical object-environment associations—whereas low-level attributes (e.g., blue) and spatial relationships (e.g., to the left) are less likely to be inferred without support from brain-derived signals. BASIC-H mitigates this confound and allows us to better isolate the contribution of genuine brain-derived information and more accurately assess the true decoding performance. BASIC-H penalizes misidentifications, such as hallucinated objects, incorrect attributes, and implausible relationships. These semantic-level discrepancies are directly reflected in the precision and recall of object categories, attributes, and relations, offering interpretable signals beyond opaque scores from pretrained networks. The segmentation-based BASIC-L scores also serve diagnostic purposes, measuring structure correspondance across four granularities: salient, semantic, instance and part. Fig. 4 presents qualitative examples illustrating how BASIC-H and BASIC-L enable fine-grained breakdowns of semantic and structural alignment. In the top reconstruction example from NeuroVLA shen2024neuro , the concepts of “dog” and “man” are correctly predicted but placed in an incorrect scene. The bottom example from MindEye2 scotti2024mindeye2 in the airplane scenario correctly predicts both semantic and structural aspects, therefore achieving a high score in both dimensions.
@@
Our evaluation follows a unified protocol across visual modalities, enabling cross-dataset dissection. The fMRI-based image reconstructions ozcelik2023brain ; scotti2023reconstructing consistently outperform EEG-based counterparts li2024visual ; zhang2025cognitioncapturer , attributable to intrinsic limitations of EEG in spatial resolution and information. The fMRI-to-video reconstructions gong2024neuroclips ; chen2023cinematic ; li2025decofuse currently struggles to preserve the structure of salient objects, but performs comparably to image modalities in terms of instance semantic identification and structural preservation. The high performance in EEG-to-video decoding liu2024eeg2video appears to reflect the simplicity of stimuli – short clips with prominent foregrounds and minimal background clutter. Future work explore more semantically rich and visually complex video scenarios. For 3D reconstruction, both fMRI-based and EEG-based visual decoding methods gao2024mind3d ; gao2024mind3d++ ; guo2025neuro3d struggle to recover even basic semantics, despite simple, canonical structures of the target object categories. The incapability in semantic identification and structure preservation highlights the need for decoding models with stronger semantic and geometric priors.
5 Conclusion
We present a unified evaluation framework for brain-based visual decoding that captures the multigranular nature of human visual perception. By integrating mask-based segmentation alignment with structured object-attribute-relation similarity, our approach enhances performance discriminability, neuroscientific validity, and semantic interpretability. The resulting BASIC metrics provide a comprehensive assessment across structural precision, inferential accuracy, and contextual coherence. We benchmark diverse brain visual decoding models across six major stimulus-neuroimaging datasets under a unified evaluation protocol, establishing the first standardized, interpretable, and extensible benchmark for this task. Our results demonstrate the framework’s broad applicability and robustness, establishing a new standard for principled evaluation in brain-guided visual decoding research.

### v2

@@ Introduction
NSD (Allen et al. 2022)
SDRecon (Takagi and Nishimoto 2023b) 9.03 13.06 38.90 49.97 15.08 21.43 17.84 1.07 4.38 6.79 11.81
BrainDiffuser (Ozcelik and VanRullen 2023) 17.96 20.21 38.98 45.85 18.66 20.78 20.09 1.94 7.86 7.82 16.65
MindEye (Scotti et al. 2023) 19.20 22.79 45.53 55.17 18.73 21.94 20.36 1.99 7.49 8.26 17.03
DREAM (Xia et al. 2024b) 23.62 26.10 46.03 57.10 21.15 24.13 21.41 2.32 9.22 8.79 19.57
MindEye2 (Scotti et al. 2024) 25.29 26.27 47.93 57.52 24.33 25.68 24.09 3.45 12.32 11.03 22.16
MindBridge (Wang et al. 2024a) 16.24 19.04 40.51 48.95 16.87 19.81 18.61 1.54 6.30 6.78 15.00
UMBRAE (Xia et al. 2024a) 21.33 24.62 40.96 48.53 18.94 21.16 20.29 2.15 8.43 8.47 17.89
NeuroPictor (Huo et al. 2024) 29.45 31.29 47.79 56.48 27.97 29.57 26.47 4.08 17.17 15.84 25.88
NeuroVLA (Shen et al. 2024) 16.03 20.85 44.09 54.42 13.84 17.54 17.57 1.34 4.38 5.57 13.54
SepBrain (Wang et al. 2024b) 21.22 23.72 45.06 53.47 20.88 23.32 21.98 2.51 8.79 8.98 18.84
UniBrain (Wang et al. 2024b) 15.24 18.02 37.48 45.09 14.99 17.75 17.45 1.06 5.54 6.07 13.79
STTM (Liu et al. 2025) 27.31 29.44 49.35 59.05 24.61 26.37 24.19 3.50 12.55 11.65 22.90
MindTuner (Gong et al. 2025) 15.38 16.66 40.74 49.02 21.46 24.50 21.49 1.97 8.16 8.13 16.98
BrainGuard (Tian et al. 2025) 25.90 28.07 49.22 58.99 23.34 25.21 23.98 3.29 10.82 10.32 21.76
EEG-Things  (Grootswagers et al. 2022)
ATM (Li et al. 2024) 13.87 18.77 39.46 50.14 22.85 34.69 22.02 2.15 11.13 17.07 17.60
CognitionCapturer (Zhang et al. 2025) 10.26 12.71 33.31 40.67 21.37 29.84 20.75 1.37 13.08 17.26 16.22
CC2017 (Wen et al. 2018)
MinD-Video (Chen, Qing, and Zhou 2023) 12.82 24.20 44.89 53.50 20.29 33.57 19.87 3.53 6.83 14.69 15.25
NeuroClips (Gong et al. 2024) 24.37 31.59 65.51 73.39 28.23 36.00 28.74 7.73 9.82 13.14 23.52
DecoFuse (Li et al. 2025) 12.32 18.07 49.87 59.46 22.33 34.63 20.87 2.74 4.06 9.80 15.31
SEED-DV (Liu et al. 2024b)
EEG2Video (Liu et al. 2024b) 27.77 32.82 57.26 69.97 22.93 31.96 23.41 3.51 3.10 7.10 20.54
fMRI-Shape (Gao et al. 2024)
MinD-3D (Gao et al. 2024) 5.69 8.10 5.69 8.10 19.91 29.35 19.38 1.45 15.96 25.30 14.72
MinD-3D++ (Gao et al. 2025) 3.80 5.23 3.80 5.23 16.03 35.83 17.96 1.83 14.94 36.19 12.62
EEG-3D (Guo et al. 2025)
Neuro-3D (Guo et al. 2025) 2.39 3.65 2.39 3.65 13.77 25.92 12.25 1.02 12.33 18.56 9.69
BASIC-H
Detailed description generation.
To support multigranular evaluation, we first prompt MLLMs to produce semantically rich descriptions from images. This description prompt extracts detailed information across key semantic components corresponding to the semantic dimensions outlined in Table 1, including scene context, objects, object attributes, inter-object relations, and camera viewpoint (when applicable). Recent studies indicate that state-of-the-art MLLMs can generate captions on par with human experts, albeit with occasional hallucinations (Lu et al. 2025; Dong et al. 2024).
Semantic graph construction.
Given detailed captions, we first segment them into individual sentences using NLTK (Loper and Bird 2004). We then extract key visual elements – objects, attributes, and relations – using a T5-based factual parser (Li et al. 2023). To handle objects mentioned across multiple sentences in the caption, we consolidate references to the same object into subcaptions and bind attributes to the correct instances, avoiding indiscriminate merging. For directional relations (e.g., subject-object pairs), we concatenate relevant sentences to maintain relational coherence (Lu et al. 2025; Dong et al. 2024).
Structured semantic matching.
The matching process aligns the extracted objects, attributes, and relations in three steps: (a) exact matching, which identifies direct lexical matches between visual elements from the reconstruction and the reference stimuli; (b) synonym matching, which aligns semantically equivalent terms, such as “building” and “edifice” for objects; “bright blue” and “azure” for attributes; “next to” and “beside” for relations; and (c) semantic matching, which computes cosine similarity for any remaining unmatched elements based on contextual meaning. This multi-step process ensures that all visual elements, whether directly mentioned or linguistically varied, are appropriately matched across decoded and reference images.
Finally, we compute semantic alignment scores based on the matching results. To mitigate the effects of omissions and hallucinations in MLLM outputs, we evaluate precision, recall, and F1 scores for each type of visual element (object, attribute, relation), based on the matching outcomes. The overall BASIC-H score is computed as a weighted sum of the three F1 scores, with weights of $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$ assigned to objects, attributes, and relations, respectively. Precision penalizes hallucinations by measuring the proportion of correctly matched elements among all generated ones, whereas recall captures omissions by evaluating how completely the reference content is recovered. Their combination in F1 score enables a balanced assessment of both over- and under-generation issues common in MLLM outputs.
BASIC-L
Semantic instance categorization.
We prompt MLLMs to generate structured semantic annotations from images. These annotations include object-centric identifications that specify spatial roles (e.g., foreground or background), semantic categories (e.g., “dog”, “tree”), and part-level component decompositions (e.g., “head”, “leg”). This process incorporates a multi-level decomposition of visual elements, capturing a hierarchical view of the scene components and facilitating further grounded segmentation and multigranular structural alignment.
Progressive granular segmentation.
The categories from these structured annotations are then used as prompts for referring expression comprehension methods (Ren et al. 2023; Kirillov et al. 2023), which produces segmentation masks for each identified object and component. The segmentation results are decided by the text and box thresholds, where only the highest-similarity boxes exceeding the box threshold and words with similarity scores above the text threshold are considered as predicted labels (Liu et al. 2024a).
Hierarchical score calculation.
We compare predicted masks against those from the reference image across four granularities: salient (foreground categories, F), binary (foreground and background categories, B), semantic (all distinct categories, S), instance (individual instances of all identified object categories, I), and part (subobject components, mostly salient foreground objects, P), computing intersection-over-union (IoU) and average precision (AP).
We aggregate saliency, semantic, instance, and part-level segmentation scores into the overall BASIC-L metric via a weighted sum of respective IoUs using weights of $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\beta_{4}$ for each granularity. This scheme prioritizes global layout and object-level coherence while accounting for fine-grained details, reflecting the limited granularity of information captured during the brain data acquisition experiments.
Experiments
Experimental Setup
Figure 2: BASIC performance.
@@
We use LLaVA-1.6-13B (Liu et al. 2023) as the default MLLM for detailed description generation and semantic instance categorization. Ground truth captions are obtained from human experts through error correction, missing element addition, and hallucination removal, based on GPT-4o-generated captions, for better alignment with human judgment. For multigranular segmentation, we employ Grounded-SAM2 (Ren et al. 2023), an open-source referring expression comprehension tool. The default weights for BASIC-H and BASIC-L are set to 4:4:2 and 3:2.5:2.5:2, respectively, reflecting the empirically informed importance of each sub-indicator in capturing semantic and structural alignment. Experiments are conducted on an NVIDIA A100 GPU. To ensure broad applicability and methodological consistency, we refrain from using MLLMs or segmentation methods specifically designed for video or 3D data. Instead, we retain an image-based framework with automatic selection of representative video frames and rendered 3D views that best capture key semantics and structure from stimuli and decoded reconstructions.
Experimental Results
While previous metrics suffer from score saturation and fail to effectively distinguish between method performances (see supplementary material), our metrics enable more nuanced evaluation. They offer clearer differentiation across objects, attributes, and relations for high-level semantics, as well as across salient, semantic, instance, and part-level granularity for low-level structure. Detailed descriptions of each dimension are provided in Table 2 and Table 3.
@@
Table 2 presents a comparison of object, attribute, and relation prediction performance across several datasets. While prior metrics often produce saturated or undifferentiated scores across models, our proposed multigranular evaluation reveals meaningful variations in object coherence, attribute accuracy, and relational plausibility, offering a more fine-grained diagnostic lens. Here, for methods on NSD (Allen et al. 2022), NeuroVLA (Shen et al. 2024), DREAM (Xia et al. 2024b), and STTM (Liu et al. 2025) achieve the highest BASIC-H scores, reflecting their strength in modeling rich visual semantics. This likely stems from detailed caption generation, complex visual semantics modeling, cross-subject training, or their combinations. A closer comparison between NeuroVLA and DREAM shows that while both models reconstruct correct instances (achieving higher precision) across the three semantic dimensions, NeuroVLA misses fewer relevant instances (better recall). In contrast, SDRecon (Takagi and Nishimoto 2023b), BrainDiffuser (Ozcelik and VanRullen 2023), and UniBrain (Wang et al. 2024b) lag behind with BASIC-H scores. These methods struggle notably with the identification of object categories, attributes, and relations, indicating limitations in handling semantics or generalizing across visual categories.
@@
Table 3 reports performance across salient, binary, semantic, instance, and part segmentation matching scores. For decoding methods on NSD, NeuroPictor (Huo et al. 2024) leads with the highest BASIC-L score, driven by its superior performance in instance and part segmentation. This reflects its ability to reconstruct fine-grained visual details and delineate object boundaries from neural signals. STTM (Liu et al. 2025) and MindEye2 (Scotti et al. 2024) perform competitively. SDRecon (Takagi and Nishimoto 2023b) and NeuroVLA (Shen et al. 2024) show relatively poor spatial structural fidelity. BrainGuard (Tian et al. 2025) achieves competitive scores in salient and semantic categories but performs less satisfactorily in instance and part segmentation. This indicates that while the model captures the overall structure of salient objects and categories, it faces challenges in distinguishing between multiple instances within the same category and identifying sub-component relationships.
Fig. 2 provides performance of visual decoding models on NSD in terms of structural alignment (BASIC-L) and semantic alignment (BASIC-H). Each point represents a model, revealing trade-offs between spatial fidelity and semantic coherence. Models closer to the top-right demonstrate stronger performance across both dimensions. For structural alignment, NeuroPictor (Huo et al. 2024), STTM (Liu et al. 2025), and MindEye2 (Scotti et al. 2024) achieve the highest BASIC-L scores, reflecting better spatial reconstruction. In contrast, semantic alignment (BASIC-H) is best captured by NeuroVLA (Shen et al. 2024), DREAM (Xia et al. 2024b), and STTM (Liu et al. 2025), indicating superior semantic preservation.
Figure 3: BASIC demonstrates stable and consistent performance in method evaluation across variations in (a) MLLMs (Liu et al. 2023), (b) prompting strategies, and (c) thresholds for box and text (Ren et al. 2023).
Discussion
@@
Open: Our evaluation pipeline is designed to be model-agnostic, avoiding reliance on proprietary or task-specific components to ensure broad applicability and reproducibility. While models like GPT-4o are shown to have stronger multimodal performance in captioning benchmarks (Lu et al. 2025; Dong et al. 2024), they are not suitable for open, large-scale benchmarking due to API restrictions and cost. We use LLaVA-1.6-13B, which balances captioning accuracy and computational efficiency. Stable: The methods under our metrics demonstrate stable and consistent performance despite variations in (a) MLLMs, (b) prompting strategies, and (c) thresholds for box and text, as shown in Fig. 3. The relative ranking and discriminative power across decoding methods remain consistent, although absolute scores may vary slightly. The stable BASIC-H results across different MLLMs and prompts indicate that our metric reliably captures semantic performance differences among methods. Similarly, the BASIC-L scores provide consistent structural evaluation across varying text and box threshold settings, where only the highest similarity boxes exceeding the box threshold and words with similarity scores above the text threshold are considered predicted labels. Versatile: For video and 3D data evaluation, instead of using modality-specific tools, we maintain the same image-based pipeline, with automatic selection of representative video frames and 3D-rendered views to ensure compatibility across formats without compromising evaluation quality.
@@
This multigranular, interpretable feedback provides fine-grained diagnostic insights into brain-based visual decoding and can help uncover blind spots. Decoding methods that rely on pretrained models benefit from powerful generative capabilities but are also susceptible to systematic biases, often introducing hallucinated details such as prototypical co-occurrence patterns. For instance, savannah may be infered when the decoded scene label is giraffe even though the original stimulus was a zoo. These hallucinations tend to affect high-level contextual information—such as scene types or typical object-environment associations—whereas low-level attributes (e.g., blue) and spatial relationships (e.g., to the left) are less likely to be inferred without support from brain-derived signals. BASIC-H mitigates this confound and allows us to better isolate the contribution of genuine brain-derived information and more accurately assess the true decoding performance. BASIC-H penalizes misidentifications, such as hallucinated objects, incorrect attributes, and implausible relationships. These semantic-level discrepancies are directly reflected in the precision and recall of object categories, attributes, and relations, offering interpretable signals beyond opaque scores from pretrained networks. The segmentation-based BASIC-L scores also serve diagnostic purposes, measuring structure correspondance across four granularities: salient, semantic, instance and part. Fig. 4 presents qualitative examples illustrating how BASIC-H and BASIC-L enable fine-grained breakdowns of semantic and structural alignment. In the top reconstruction example from NeuroVLA (Shen et al. 2024), the concepts of “dog” and “man” are correctly predicted but placed in an incorrect scene. The bottom example from MindEye2 in the airplane scenario correctly predicts both semantic and structural aspects, therefore achieving a high score at both dimensions.
@@
Our evaluation follows a unified protocol across visual modalities, enabling cross-dataset dissection. The fMRI-based image reconstructions (Ozcelik and VanRullen 2023; Scotti et al. 2023) consistently outperform EEG-based counterparts (Li et al. 2024; Zhang et al. 2025), attributable to intrinsic limitations of EEG in spatial resolution and information. The fMRI-to-video reconstructions (Gong et al. 2024; Chen, Qing, and Zhou 2023; Li et al. 2025) currently struggles to preserve the structure of salient objects, but performs comparably to image modalities in terms of instance semantic identification and structural preservation. The high performance in EEG-to-video decoding (Liu et al. 2024b) appears to reflect the simplicity of stimuli – short clips with prominent foregrounds and minimal background clutter. Future work explore more semantically rich and visually complex video scenarios. For 3D reconstruction, both fMRI-based and EEG-based visual decoding methods (Gao et al. 2024, 2025; Guo et al. 2025) struggle to recover even basic semantics, despite simple, canonical structures of the target object categories. The incapability in semantic identification and structure preservation highlights the need for decoding models with stronger semantic and geometric priors.
Conclusion
We present a brain-based visual decoding evaluation framework that captures the multigranular nature of human visual perception. By integrating mask-based segmentation alignment with structured object-attribute-relation similarity, our approach enhances performance discriminability, neuroscientific validity, and semantic interpretability. The resulting BASIC metrics provide a comprehensive assessment across structural precision, inferential accuracy, and contextual coherence. We benchmark diverse brain visual decoding models across six major stimulus-neuroimaging datasets under a unified evaluation protocol, establishing the first standardized, interpretable, and extensible benchmark for this task.
Acknowledgements. This work was supported by a UKRI Future Leaders Fellowship [grant number G104084].
Appendix
Supplemental Material for “Multigranular Evaluation for Brain Visual Decoding”